Louise is four times as old as Maria. However, 6 years from now she will be only three times as old as Maria. Find their ages.
Answer by robertb(5830)

L = 4M.
L+6 = 3(M+6).
By direct substitution,
4M +6 = 3M+18,
M = 12. So Maria is 12 years old. Louise is 48 years old.